Molecular formula c5h10o can hav structural isomers for aldehydes. draw all of them.

Grade:12

1 Answers

JAYAVARDHAN BAIRABATHINA
10 years ago

1)  CH3-C-CH2-CH2-CH3   2-pentanone
                I I
                O

2) CH3-CH2-C-CH2-CH3     3-pentanone
                        I I
                        O

3) CH3-C-CH-CH3               Methyl isopropyl ketone
              I I   I
              O   CH3

1'')  CH3-CH2-CH2-CH2-HC=O  pentanal

2'')  CH3-CH2-O-CH=CH-CH3